How many atoms of gold are there in 7g of a karat gold ring?


1
Expert's answer
2022-02-28T11:05:51-0500

Solution:

The molar mass of gold (Au) is 196.97 g/mol

Therefore,

Moles of Au = (7 g Au) × (1 mol Au / 196.97 g Au) = 0.03554 mol Au


One mole of any substance contains 6.022×1023 particles.

Therefore,

Number of Au atoms = (0.03554 mol Au) × (6.022×1023 Au atoms / 1 mol Au) = 2.14×1022 Au atoms


Answer: 2.14×1022 atoms of gold